# Physics Context Builders: A Modular Framework for Physical Reasoning in Vision-Language Models

### v2

@@ 3.3 Ablation Studies: Not All Data is Equal
Table 3: Effect of including different QA types in fine-tuning PaliGemma-3B on Falling Tower.
Fine-tuning Data Descriptive [acc.] Stability [acc.] num. obj. num. $\uparrow\downarrow↑ ↓ obj. absent\uparrow\downarrow$↑ ↓ obj. stable tower stable No Fine-tuning 50.9 21.2 8.5 51.0 39.1 Stability QAs 52.4 26.0 41.2 84.4 86.5 Descriptive QAs 100.0 100.0 100.0 51.0 39.1 All QAs 100.0 100.0 100.0 84.6 87.6
@@
Specificity of QA types in fine-tuning.
Table 3 shows the impact of using different subsets of QA pairs for fine-tuning on Falling Tower. Fine-tuning with only descriptive questions achieves perfect descriptive accuracy but fails to improve stability prediction (remaining at baseline levels of 51.0% and 39.1%). Interestingly, fine-tuning with only stability questions not only improves stability prediction substantially but also yields moderate improvements in descriptive tasks, particularly for object identification (41.2% vs. 8.5% baseline). However, the best performance comes from fine-tuning with the complete QA dataset. This highlights the need for carefully selecting the fine-tuning data to cover all potential QA types, which can be seen as a limitation of targeted fine-tuning in VLMs.
@@
[…]
@@ 4 Physics Context Builders: A Modular Approach to Physical Reasoning
While fine-tuning with physics-focused data can significantly improve physical reasoning in VLMs, this approach faces practical limitations. Fine-tuning state-of-the-art foundation models like GPT-4o or Gemini is often expensive or even impossible due to their closed-source nature and computational requirements. Furthermore, the task-specific nature of fine-tuning, as seen in section 3, means separate models may be needed for different physical reasoning tasks, which limits the generalizability of the approach. To address these challenges, we introduce Physics Context Builders (PCBs) - a modular and efficient approach that enhances physical reasoning capabilities of existing VLMs without modifying them directly. PCBs leverage the strong in-context learning abilities of large language models (Brown,, 2020; Wei et al.,, 2022), which have shown impressive performance with textual scene descriptions (Nagar et al.,, 2024).
@@ 4.1 Methodology
PCBs are specialized VLMs fine-tuned on simulation data to generate detailed physical descriptions of visual scenes. Rather than directly answering questions, PCBs act as perception modules that translate visual inputs into rich textual descriptions capturing the physical properties and dynamics of a scene. These descriptions then serve as enhanced context for larger VLMs, enabling more accurate physical reasoning through in-context learning. Figure 1 provides an overview of our approach. PCBs offer several advantages:
@@ 4.1.1 Physical Context Generation
Using annotations from the simulator, PCBs generate question-agnostic descriptions that capture the physical essence of a scene. We consider two approaches for context generation: (1) Human-like Narration (HN) that uses natural language descriptions of the scene’s physical properties and events that can be more intuitive for foundation models, and (2) Structured Physics (SP), which provides frame-by-frame structured observations in a format similar to a physics simulation output, with standardized tags for physical properties. This structured approach helps models capture precise temporal relationships and physical properties. For the stability detection task in Falling Tower, we generate descriptions, following the human-like narration, including object types, locations, and stacking arrangements. Figure 4 provides examples for these description formats.
Figure 4: PCB generation formats. (left) Human-like narration for CLEVRER. (top right) Structured physics for CLEVRER. (bottom right) Human-like narration for Falling Tower.
@@ 4.1.2 Training PCBs
    Similar to section 3, we apply LoRA-based fine-tuning (Hu et al.,, 2021), minimizing the auto-regressive negative log-likelihood of the context descriptions conditioned on the input video/image. See section A.2 for more details.
@@
    The fine-tuning process targets both vision and language components, ensuring proper alignment between visual features and linguistic representations.
@@ 4.2 Experimental Results
Table 4: Performance of VLMs augmented with Physics Context Builders (PCBs) on the Falling Tower benchmark. HN refers to Human Narration style PCB. The numbers in parentheses indicate improvements over the respective zero-shot baselines. The second number after the slash is the Sim2Real accuracy and the third number after the second slash is the F1 score on Sim2Real.
Category Model Descriptive [sim acc. / real acc.] Stability [sim acc. / real acc. / real F1] num. obj. num. $\uparrow\downarrow↑ ↓ obj. absent\uparrow\downarrow$↑ ↓ obj. stable tower stable Zero-shot CoT GPT-4o 99.3 / 100.0 91.4 / 100.0 99.4 / 95.0 56.9 / 60.0 / 63.3 59.6 / 55.0 / 54.2 GPT-4o-mini 94.9 / 89.5 61.3 / 84.2 87.9 / 73.7 49.0 / 52.6 / 55.0 53.1 / 36.8 / 19.8 Gemini 1.5 Pro 97.2 / 95.0 89.9 / 100.0 97.8 / 95.0 54.6 / 80.0 / 80.0 60.5 / 60.0 / 60.1 VLM + PCB (HN) GPT-4o-PCB 99.5 / 100.0 97.6 / 100.0 99.5 / 95.0 76.7 / 75.0 / 73.8 85.1 / 70.0 / 65.6 (+0.2) / (0.0) (+6.2) / (0.0) (+0.1) / (0.0) (+19.8) / (+15.0) / (+10.5) (+25.5) / (+15.0) / (+11.4) GPT-4o-mini-PCB 99.9 / 95.0 74.3 / 90.0 97.5 / 95.0 75.0 / 70.0 / 73.0 84.7 / 40.0 / 33.8 (+5.0) / (+5.5) (+13.0) / (+5.8) (+9.6) / (+21.3) (+26.0) / (+17.4) / (+18.0) (+31.6) / (+3.2) / (+14.0) Gemini 1.5 Pro-PCB 97.9 / 100.0 97.5 / 100.0 97.4 / 94.7 75.9 / 73.7 / 76.7 84.9 / 57.9 / 59.1 (+0.7) / (+5.0) (+7.6) / (0.0) (-0.4) / (-0.3) (+21.3) / (-6.3) / (-3.3) (+24.4) / (-2.1) / (-1.0)
Table 5: Performance of VLMs augmented with Physics Context Builders (PCBs) on the CLEVRER benchmark. HN is Human Narration and SP is Structured Physics. The numbers in parentheses indicate improvements over the respective zero-shot baselines.
@@ 4.2.1 PCB Performance on Falling Tower
Table 4 shows the performance of foundation models augmented with PCBs on the Falling Tower benchmark:
@@ 4.2.2 PCB Performance on CLEVRER
Table 5 presents results for the more challenging CLEVRER benchmark, which requires reasoning about dynamic physics in videos:
@@
  • Limited gains in counterfactual reasoning: While PCBs improve counterfactual reasoning, the gains are more modest (around 1.7-9.5%), highlighting the inherent difficulty of this task even with enhanced context.
Remark. For CLEVRER, we omit PCB evaluation on predictive questions since our current PCBs are designed to describe observed scenes rather than predict future events. This limitation could be addressed in future work by developing PCBs that can describe potential future states, which can be easily obtained using simulation.
@@ 4.3 Multi-Agent Framework for PCB Integration
Figure 5: Multi-agent framework for PCB integration. A triage agent selects the appropriate PCB based on the user query, and the PCB generates a detailed scene description that enriches the context of the foundation model.
Table 6: Performance metrics for selecting the correct context builder agent in a multi-agent framework (two agents).
Given the ability of foundational models to interpret the overall context of scenes effectively, we explore whether they can reliably select the appropriate PCB when provided with a question-image pair. We utilize a multi-agent triage model inspired by OpenAI’s Swarm architecture (Bigio et al.,, 2024). As illustrated in Figure 5, our multi-agent framework consists of:
@@
In our evaluation, each input consists of a natural language question from a QA dataset paired with a corresponding scene image. This input is initially processed by GPT-4o or GPT-4o-mini that routes the query to one of two specialized PCBs: the PCB for the Falling Tower dataset, or PCB designed for analyzing motion and object interactions as in the CLEVRER dataset. As shown in Table 6, both GPT-4o and GPT-4o-mini achieve excellent F1-Scores, effectively routing queries to the appropriate PCB. These results suggest that PCBs in a multi-agent framework offer a promising approach, with foundation models capable of reliably selecting the correct PCB based on the question-image pair.
@@ 5 Discussion
Our results demonstrate that Physics Context Builders (PCBs) offer a promising approach to enhancing physical reasoning in Vision-Language Models. PCBs increase the average performance of GPT-4o (GPT-4o-mini) by 11.1% (11.8%) on CLEVRER and 8.2% (13.8%) on Falling Tower. These improvements are particularly notable given that no modification to the foundation models was required. Despite these gains, large VLMs still exhibit sub-optimal performance in tasks requiring deeper reasoning, such as counterfactual questions (with improvements limited to 1.7-9.5%) and explanatory questions; while enhancing visual perception through PCBs is important, especially for descriptive tasks, the subpar performance on more involved reasoning tasks calls for additional interventions to reach a comprehensive physical understanding.
Besides the performance gain, our findings confirm the value of simulation data in addressing the limitations of VLMs. Unlike simulation-in-the-loop approaches (Wu et al.,, 2015, 2017; Liu et al.,, 2022; Ding et al., 2021b, ) that require computationally expensive simulators during inference, our method uses simulation only to generate synthetic data (context descriptions) for fine-tuning smaller VLMs. This approach is more efficient at inference while still leveraging the rich annotations that simulations can provide. The effectiveness of PCBs in generalizing from simulated to real-world data further supports this approach, as demonstrated by the successful Sim2Real transfer in the Falling Tower benchmark.
PCBs provide a modular, efficient framework for enhancing the physical reasoning capabilities of foundation models without requiring direct fine-tuning. Generating rich physical context from visual inputs uplifts the perceptual ability of foundation models, resulting in more accurate reasoning across diverse physical tasks, from static stability to dynamic collision detection. The multi-agent framework further enhances this approach by enabling adaptive selection of specialized PCBs based on the specific reasoning task.
@@ 6 Limitations and Future Work
While our work demonstrates the effectiveness of PCBs across the benchmarks, several limitations remain. First, our current benchmarks focus on a relatively constrained set of physical phenomena—primarily rigid body dynamics and stability. This limits our ability to evaluate how well these approaches might generalize to the full spectrum of physical reasoning that humans perform intuitively, such as reasoning about fluid dynamics or object manipulation.
Second, our framework requires annotated data, which naturally comes from simulation. For unannotated real-world videos, such as those available on YouTube, the lack of structured annotations presents a challenge. An important open problem is how PCBs could be used to extract detailed physical descriptions from real-world videos without explicit annotations, potentially through self-supervised learning or by leveraging other foundation models to generate pseudo-annotations for real data.
Third, the performance improvements, while substantial, still leave room for further enhancement, particularly for complex reasoning tasks like counterfactual/predictive questions. Since PCBs inherently perform direct translation from visual signals to text, they result in strong descriptive capabilities of the foundation models. However, this translation process does not capture all the visual cues needed for more complex reasoning. Hence, we hypothesize that more comprehensive textual descriptions, possibly containing counterfactual/predictive information, could significantly improve the performance of PCBs on these challenging tasks.
A promising future direction is to tackle more complex physical reasoning tasks that require larger-scale simulations. This might involve integrating advanced physics engines that simulate phenomena like fluid dynamics, deformable materials, and articulated mechanisms. Notably, while these simulations could become computationally expensive for simulation-in-the-loop approaches during inference, our PCB framework uses simulation data only for fine-tuning and therefore would not add significant computational burden during inference.
Another avenue for future work is exploring how multiple PCBs could be composed or chained together to handle scenarios requiring reasoning about multiple physical phenomena simultaneously. For instance, reasoning about a scene involving both rigid body dynamics and fluid interactions might benefit from specialized PCBs for each phenomenon, with their outputs combined to provide comprehensive context to the foundation model.
@@ Appendix A Training Details
Here, we provide the training details, including the hyper-parameters, for both QA fine-tuning and PCB training tasks. All of our training was performed on Nvidia’s A100 DGX systems.
@@ A.1 QA fine-tuning
To fine-tune the PaliGemma-3B on the question-answer datasets, we apply the LoRA fine-tuning scheme by targeting the attention weights in both vision and language modules, as well as the fully-connected MLP layers, the multi-modal projector layers, the embedding tokens, the patch embedding, and the positional embedding. Table 7 (left) shows other hyper-parameters used for QA fine-tuning.
Table 7: Hyper-parameters used to fine-tune the PaliGemma-3B models (left) on the QA datasets and (right) as PCB modules.
Hyper-parameter Falling Tower CLEVRER
@@
[…]
Hyper-parameter Falling Tower CLEVRER

### v3

@@ 3.3 Ablation Studies: Not All Data is Equal
Table 3: Effect of fine-tuning PaliGemma-3B on different QA types in Falling Tower.
Fine-tuning Data Descriptive [acc.] Stability [acc.] num. obj. num. obj. $\uparrow\downarrow$ obj. $\uparrow\downarrow$ obj. stable tower stable No Fine-tuning 50.9 21.2 8.5 51.0 39.1 Stability QAs 52.4 26.0 41.2 84.4 86.5 Descriptive QAs 100.0 100.0 100.0 51.0 39.1 All QAs 100.0 100.0 100.0 84.6 87.6
Specificity of QA types in fine-tuning.
Tab. 3 shows the impact of using different subsets of QA pairs for fine-tuning on Falling Tower. Fine-tuning with only descriptive questions achieves perfect descriptive accuracy but fails to improve stability prediction (remaining at the baseline level). Interestingly, fine-tuning with only stability questions not only improves stability prediction substantially but also yields moderate improvements in descriptive tasks, particularly for object identification (41.2% vs. 8.5% baseline). However, the best performance comes from fine-tuning on the full QA dataset. This highlights the need to carefully select data to cover all QA types, which can be seen as a limitation of targeted fine-tuning in VLMs.
@@
Fig. 3 illustrates how accuracy varies with the amount of training data for different QA types in Falling Tower. Descriptive questions reach near-saturation with only $\sim$10% of the data, showing the model can quickly learn basic scene understanding. In contrast, stability questions benefit more from extra data, though with diminishing returns. Even with 100% of the available data, the model does not reach perfect performance on stability questions (87.6% for tower stability). This shows that increasing dataset size alone is insufficient; improvements in physical reasoning also require diverse, targeted supervision.
@@ 4 Physics Context Builders: A Modular Approach to Physical Reasoning
While fine-tuning with physics-focused data can significantly enhance physical reasoning in VLMs, it presents notable practical limitations. Fine-tuning state-of-the-art foundation models like GPT-4o or Gemini is often expensive or even impossible due to their closed-source nature and computational requirements. Furthermore, the task-specific nature of fine-tuning, as seen in Sec. 3, means separate models may be needed for different physical reasoning tasks, which limits the generalizability of the approach. To address these challenges, we introduce Physics Context Builders (PCBs) – a modular and efficient approach that enhances physical reasoning capabilities of existing VLMs without modifying them directly. PCBs leverage the strong in-context learning abilities of large language models [8, 41], which have shown impressive performance with textual scene descriptions [31].
@@ 4.1 Methodology
PCBs are specialized VLMs fine-tuned on simulation data to generate detailed physical descriptions of visual scenes. Rather than directly answering questions, PCBs act as perception modules that translate visual inputs into rich textual descriptions capturing the physical properties and dynamics of a scene. These descriptions then serve as enhanced context for larger VLMs, enabling more accurate physical reasoning through in-context learning. Fig. 1 provides an overview of our approach. PCBs offer several advantages:
@@ 4.1.1 Physical Context Generation
Using annotations from the simulator, PCBs generate question-agnostic descriptions that capture the physical essence of a scene. We consider two context types:
(1) Human-like Narration (HN): Produces intuitive natural language descriptions of the scene’s physical properties and events that can be more aligned with foundation models:
(2) Structured Physics (SP): Provides frame-by-frame structured observations in a format similar to a physics simulation output, with standardized tags for physical properties. This structured approach helps models capture precise temporal relationships and physical properties:
@@ 4.1.2 Training PCBs
    We apply LoRA-based fine-tuning [18], minimizing the auto-regressive negative log-likelihood of the context descriptions conditioned on the input video/image. See Sec. A.2 for more details.
@@
    Fine-tuning jointly optimizes both vision and language components to ensure alignment between visual features and linguistic representations.
@@ 4.2 Experimental Results
Table 4: Performance of foundation models augmented with Physics Context Builders (PCBs) on the Falling Tower benchmark. HN refers to Human Narration style PCB. The numbers in parentheses indicate improvements over the respective zero-shot baselines. The second number after the slash is the Sim2Real accuracy and the third number after the second slash is the F1 score on Sim2Real.
Category Model Descriptive [sim acc. / real acc.] Stability [sim acc. / real acc. / real F1] num. obj. num. obj. $\uparrow\downarrow$ obj. $\uparrow\downarrow$ obj. stable tower stable Zero-shot CoT GPT-4o 99.3 / 100.0 91.4 / 100.0 99.4 / 95.0 56.9 / 60.0 / 63.3 59.6 / 55.0 / 54.2 GPT-4o-mini 94.9 / 89.5 61.3 / 84.2 87.9 / 73.7 49.0 / 52.6 / 55.0 53.1 / 36.8 / 19.8 Gemini 1.5 Pro 97.2 / 95.0 89.9 / 100.0 97.8 / 95.0 54.6 / 80.0 / 80.0 60.5 / 60.0 / 60.1 VLM + PCB (HN) GPT-4o-PCB 99.5 / 100.0 97.6 / 100.0 99.5 / 95.0 76.7 / 75.0 / 73.8 85.1 / 70.0 / 65.6 (+0.2) / (0.0) (+6.2) / (0.0) (+0.1) / (0.0) (+19.8) / (+15.0) / (+10.5) (+25.5) / (+15.0) / (+11.4) GPT-4o-mini-PCB 99.9 / 95.0 74.3 / 90.0 97.5 / 95.0 75.0 / 70.0 / 73.0 84.7 / 40.0 / 33.8 (+5.0) / (+5.5) (+13.0) / (+5.8) (+9.6) / (+21.3) (+26.0) / (+17.4) / (+18.0) (+31.6) / (+3.2) / (+14.0) Gemini 1.5 Pro-PCB 97.9 / 100.0 97.5 / 100.0 97.4 / 94.7 75.9 / 73.7 / 76.7 84.9 / 57.9 / 59.1 (+0.7) / (+5.0) (+7.6) / (0.0) (-0.4) / (-0.3) (+21.3) / (-6.3) / (-3.3) (+24.4) / (-2.1) / (-1.0)
Table 5: Performance of foundation models augmented with Physics Context Builders (PCBs) on the CLEVRER benchmark. HN is Human Narration and SP is Structured Physics. The numbers in parentheses indicate improvements over the respective zero-shot baselines.
@@ 4.2.1 PCB Performance on Falling Tower
Tab. 4 shows the performance of foundation models augmented with PCBs on Falling Tower. For additional experiments, see Sec. D.2. Several key findings:
@@ 4.2.2 PCB Performance on CLEVRER
Tab. 5 presents results for CLEVRER, a more challenging benchmark that requires dynamic physical reasoning:
@@
  • Limited gains in counterfactual reasoning: While PCBs improve counterfactual reasoning (1.7–9.5% gains), the improvements are more modest, reflecting the intrinsic complexity of this task, even with enriched context.
Remark. For CLEVRER, we omit PCB evaluation on predictive questions since our current PCBs are designed to describe observed scenes rather than predict future events. Future work could address this by developing predictive PCBs that generate plausible future scene descriptions based on simulation rollouts.
@@ 4.3 Multi-Agent Framework for PCB Integration
Figure 4: Multi-agent framework for PCB integration. A triage agent selects the appropriate PCB based on the user query, and the PCB generates a detailed scene description that enriches the foundation model’s context.
Given the ability of foundational models to interpret the overall context of scenes effectively, we explore whether they can reliably select the appropriate PCB when provided with a question-image pair. We utilize a multi-agent triage model inspired by OpenAI’s Swarm architecture [7]. As illustrated in Fig. 4, our multi-agent framework consists of:
@@
In our evaluation, each input consists of a natural language question from a QA dataset paired with a corresponding scene image. This input is initially processed by GPT-4o or GPT-4o-mini that routes the query to one of two specialized PCBs: the PCB for the Falling Tower dataset, or PCB designed for analyzing motion and object interactions as in the CLEVRER dataset. As shown in Tab. 6, both GPT-4o and GPT-4o-mini achieve excellent F1-Scores, effectively routing queries to the appropriate PCB. These results suggest that PCBs in a multi-agent framework offer a promising approach, with foundation models capable of reliably selecting the correct PCB based on the question-image pair.
Table 6: Accuracy and F1-scores for selecting the correct Physics Context Builder using the triage agent in a two-agent setup.
@@ 5 Discussion
Our results demonstrate that Physics Context Builders (PCBs) offer a promising approach to enhancing physical reasoning in Vision-Language Models (VLMs). PCBs increase the average performance of GPT-4o (GPT-4o-mini) by 11.1% (11.8%) on CLEVRER and 8.2% (13.8%) on Falling Tower. These improvements are particularly notable given that no modification to the foundation models was required. Despite these gains, large VLMs still exhibit sub-optimal performance in tasks requiring deeper reasoning, such as counterfactual questions (with improvements limited to 1.7-9.5%) and explanatory questions. While enhancing visual perception through PCBs is important, especially for descriptive tasks, the subpar performance on more involved reasoning tasks calls for additional interventions to achieve a comprehensive physical understanding.
Besides the performance gain, our findings confirm the value of simulation data in addressing the limitations of VLMs. Unlike simulation-in-the-loop approaches [42, 43, 29, 12], which require computationally expensive simulators during inference, our method uses simulation only to generate synthetic data (context descriptions) for fine-tuning smaller VLMs. This approach is more efficient at inference while still leveraging the rich annotations that simulations can provide. The effectiveness of PCBs in generalizing from simulated to real-world data further supports this approach, as demonstrated by the successful Sim2Real transfer in the Falling Tower benchmark.
PCBs provide a modular and efficient framework for enhancing the physical reasoning capabilities of foundation models without requiring direct fine-tuning. Generating rich physical context from visual inputs enhances the perceptual ability of foundation models, resulting in more accurate reasoning across diverse physical tasks, from static stability to dynamic collision detection. The multi-agent framework further enhances this approach by enabling adaptive selection of specialized PCBs based on the specific reasoning task.
@@ 6 Limitations and Future Work
While our work demonstrates the effectiveness of PCBs across benchmarks, several limitations remain. First, our current benchmarks focus on a relatively constrained set of physical phenomena—primarily rigid body dynamics and stability. This limits our ability to evaluate how well these approaches generalize to the full spectrum of physical reasoning that humans perform intuitively, such as fluid dynamics or object manipulation.
Second, our framework requires annotated data, which naturally comes from simulation. For unannotated videos, such as those available on YouTube, the lack of structured annotations presents a challenge. An important open problem is how PCBs could extract detailed physical descriptions from real-world videos without explicit annotations, potentially through self-supervised learning or by leveraging other foundation models to generate pseudo-annotations.
Third, the performance improvements, while substantial, still leave room for further enhancement, particularly for complex reasoning tasks like counterfactual/predictive questions. Since PCBs inherently perform direct translation from visual signals to text, they lead to strong descriptive capabilities in the foundation models. However, this translation process does not capture all the visual cues needed for more complex reasoning. Hence, we hypothesize that more comprehensive textual descriptions, possibly containing counterfactual/predictive information, could significantly improve the performance of PCBs on these challenging tasks.
A promising future direction is to tackle more complex physical reasoning tasks that require larger-scale simulations. This might involve integrating advanced physics engines that simulate phenomena like fluid dynamics, deformable materials, and articulated mechanisms. Notably, while these simulations could become computationally expensive for simulation-in-the-loop approaches during inference, our PCB framework uses simulation data only for fine-tuning and therefore does not add significant computational burden at inference time.
Another avenue for future work is to explore how multiple PCBs could be composed or chained together to handle scenarios requiring reasoning about multiple physical phenomena simultaneously. For instance, reasoning about a scene involving both rigid body dynamics and fluid interactions might benefit from specialized PCBs for each phenomenon, with their outputs combined to provide comprehensive context to the foundation model.
Acknowledgements
We thank Brian Jeong for 3D printing the parts used in the Falling Tower dataset. Resources for this research were primarily provided by Autodesk Research, with partial support from the Province of Ontario, the Government of Canada through CIFAR, and sponsors of the Vector Institute.
@@ Appendix A Training Details
Here, we provide the training details, including the hyperparameters, for both QA fine-tuning and PCB training tasks. All training was performed on NVIDIA A100 DGX systems.
@@ A.1 QA fine-tuning
To fine-tune PaliGemma-3B on the question-answer datasets, we apply the LoRA fine-tuning scheme by targeting the attention weights in both the vision and language modules, as well as the fully connected MLP layers, multi-modal projector layers, embedding tokens, patch embedding, and positional embedding. Tab. 7 (left) shows other hyperparameters used for QA fine-tuning.
Table 7: Hyperparameters used to fine-tune the PaliGemma-3B models on the QA datasets (left) and as PCB modules (right).
Hyperparameter Falling Tower CLEVRER
@@
Compute time $\sim$ 3.5 hours $\sim$ 37 hours
Hyperparameter Falling Tower CLEVRER